OpenStudy (ms_sp):

1. How many ounces of Asian treasures tea that sells for $7.50/oz should be mixed with pearadise tea that sells for $5.00/oz so that a 60-oz mixture is obtained that will sell for $6.00/oz? 2. How many milliliters of a 15% acid solution and how many millimeters of a 3% acid solution must be mixed to get a 45 mL of 7% alcohol solution?

OpenStudy (ms_sp):

@AlexaTiger

OpenStudy (ms_sp):

@triciaal

OpenStudy (ms_sp):

@brokenstanzas

OpenStudy (alexatiger):

Do you have a formula or a way to work this? //Or do you have a guess?

OpenStudy (brokenstanzas):

I think this would be solved by using ratios but I'm not too sure

OpenStudy (ms_sp):

Don't have a formula, all I know is x+y=45. @AlexaTiger And, @brokenstanzas : We don't do ratios.

OpenStudy (alexatiger):

Actually, I'm pretty sure the way I was taught was with ratios. Cool, @brokenstanzas We were looking at number one when I said that :) No ratios for number one either?

OpenStudy (ms_sp):

None @AlexaTiger

OpenStudy (alexatiger):

Alrighty, thanks. I can't figure this out, I'd guess at the first one and I haven't done Chemistry before. Sorry :(

OpenStudy (brokenstanzas):

Maybe \[7.50x+5.00y=6.00(60)\] I'm kinda guessing at this point

OpenStudy (ms_sp):

It's not chemistry, it's math you guys!